One positive number is 8 times another positive number. The difference between the two numbers is 126 . Find the two numbers
Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

One positive number is 8 times another positive number. The difference between the two numbers is 126 . Find the two numbers
-------------
The difference is 7 times the number.
126/7 = 18
--> 18 and 144

Answer by addingup(3677)

x = 8y Since x= 8y, we will use this value for x next.
x-y = 126 Substitute for x:
8y-y = 126
7y = 126
y = 18
------------------------------------
Check:
x = 8y
x = 8(18)
x = 144
and
144 - 18 = 126 We have the correct answer.
J